Week 3: Jan. 21-25-20 QUESTION7 Assume that the market for Good X is defined as follows: QD- 64-16P and QS O$24.50 16P-8. If the government imposes a price floor in this market at $3.00, what will consumer surplus be? $49.00 $9.00 $8.00 $32.00 : : QUESTION8 Assume that the market for Good X is defined as follows: QD- 64-16P and QS 16P-8. If the government imposes a price floor in this market at $3.00, what will producer surplds be $24.00 $49.00 O $55.00 O$32.00 QUESTION9 quim e that the r arket for Good Xi defined as follo s OD·64-16P and economy S-16p.8rthe government imposes a pr ef oor in this market at S3 00, what will be the total loss in welfare to the $49.00 $9.00 $8.00 ○There is no loss
Answer #1

7. Qd= 64-16P

When Qd=0 ,P =4

Qs= 16P-8

When Qs=0 ,P =0.5

At the price floor of $3 .

Qd= 64-16(3) = 16

Consumer surplus is the triangle area above $3 and below the demand curve.

Consumer surplus = (0.5)(4-3)(16) = $8

Hence, option (D) is correct.

8. When Q=16 , it touches the supply curve at $1.5 .

Producer surplus is the area above the supply curve and below the price $3. It is the sum of rectangle and triangle area .

Producer surplus = (3-1.5)(16) +(0.5)(1.5-0.5)(16)

= 24+ 8

= $32

Hence,option( E) is correct.

9. At equilibrium Qd= Qs

64-16P = 16P-8

32P = 72

P= $2.25 (Equilibrium price)

Q= 64-16(2.25) = 28 (Equilibrium quantity)

Total loss in welfare is shown by the area of the sum of two triangles .

= (0.5)(3-2.25)(28-16) +(0.5)(2.25-1.5)(28-16)

= 4.5 + 4.5

= $9

Hence, option (C) is correct.

3 し 15-28